# Fuzzy Dark Matter Constraints from the Hubble Frontier Fields

### v1

@@ 4 Sensitivity to Model Assumptions
The faint end of the UVLF, even including the faint, lensed galaxies of B22, is well-fit by the double power-law luminosity-halo mass relation (Sipple & Lidz, 2024). This assumption is also consistent with results from zoom-in hydrodynamic simulations of galaxy formation from the FIRE collaboration (Ma et al., 2018). Yet, it also may be the case that feedback effects (which arrest star formation) are weaker (e.g. if star formation is exceptionally “bursty” Furlanetto & Mirocha 2022; Sun et al. 2023) at the faint-end, or non-hierarchical structure formation in FDM may lead to younger, brighter stars in the lowest-mass halos (Corasaniti et al., 2017; Ni et al., 2019; Shen et al., 2023). [^7] The possibility of stronger feedback effects in the smallest halos (e.g. from radiative feedback during the reionization process, Borrow et al. 2023) is also not captured by this simple model. However, we do not find a clear turnover in the current UVLF data and, unlike the previous scenarios, neglecting this effect in our analyses leads to only a more conservative bound on FDM mass.
@@ 4.1 Cumulative UVLF
The strong suppression in the FDM halo mass function implies the cumulative number density of halos asymptotes to some finite value. This can be compared with the total observed abundance of galaxies above the limiting luminosity of the B22 data. Assuming only that there are one or fewer galaxies occupying each halo on average, which is expected theoretically at the faint end (Ma et al., 2018) and supported empirically by stringent bounds on satellite fractions at high redshift (Harikane et al., 2022), this provides an alternate, more conservative bound on FDM. While our previous approach uses the full shape of the UVLF and delivers tighter bounds, the cumulative UVLF technique involves fewer assumptions regarding how galaxies populate their host halos. Although this approach has been applied in previous work (Schive et al., 2016; Menci et al., 2017; Ni et al., 2019), we emphasize here its robustness to galaxy-halo connection modeling uncertainties.
@@
Given the requirement that fewer than one galaxy resides in each halo, on average, and with $M_{\rm min}$ denoting the minimum mass of dark matter halos hosting star formation, the halo mass function is inconsistent with the cumulative galaxy number density implied by UVLF measurements unless:
@@ 4.2 Modelling Enhanced Star Formation in Small Halos
If star formation proceeds differently in the smallest halos, we may still be able to model this alternate behavior as a simple additional power law in our UVLF modeling procedure. To this end, we introduce two additional degrees of freedom in our $L_{c}(M)$ relation, a new power law index, $p_{2}$, and its relevant mass scale, $M_{2}$, (similar to the “shallow-slope” model of Sipple & Lidz 2024) and refit our parameters.
Sampling these seven parameters with MCMC, we find this model produces a similar constraint to the cumulative UVLF, (here $m_{22}>5.8$ at $2\sigma$) (Table 2). If the very smallest halos are actually bright enough to be included in B22 (necessitating 23$\sim$ 2-3$\times$ 2 - 3 × the predicted star formation efficiency of our fiducial model), this could somewhat compensate for the fewer halos at a given mass in FDM. While this may indicate more efficient star formation at the faint-end, including additional parameters always presents a danger of overfitting. This new behavior coincidentally onsets near the edge of observational capabilities, where the errors bars are relatively large. Further progress may require improving the statistical precision of the faint-end UVLF measurements and additional simulations exploring star formation in small mass dark matter halos.
@@ 4.3 Effect of Prior
Another common choice in the literature is a flat prior in $\log_{10}(m_{\rm FDM})$ (Rogers & Peiris, 2021; Dentler et al., 2022; Winch et al., 2024), or in the log of the suppression halo mass scale (Nadler et al., 2021; Banik et al., 2021). This gives equal weight to different regions in log-space, but only over a finite interval, and so this cannot explicitly consider CDM where $m_{22}\rightarrow\infty$. The upper bounds on these intervals are chosen taking into account experimental sensitivity, although if the likelihood is not well-peaked in the prior region this choice can somewhat influence the confidence intervals on $m_{22}$. With a prior of $\log_{10}(m_{22})\in[-1,2]$ (instead of our fiducial case of $1/m_{22}\in[0,2]$) our $2\sigma$ UVLF modeling constraint shifts only slightly from $m_{22}>15$ to $m_{22}>18$ (Tables 1,2).
@@ 5 FDM Halo Mass Function Modelling Uncertainties
Although a great deal of progress has been made recently, there is not yet a single simulation which spans the large dynamic range in scale need to resolve FDM effects near the local de Broglie wavelength, $\lesssim 100$ pc, while capturing statistical behavior on cosmological scales, $\gtrsim 100$ Mpc (Zhang et al., 2018; Li et al., 2019; Hui, 2021; May & Springel, 2023; Laguë et al., 2024). The halo mass function relation of Equation 1 is only one approximation for how structure formation occurs in FDM. As mentioned in Section 2.1, it is fit to $N$-body simulations of Schive et al. (2016) with FDM initial conditions (IC) (including a suppressed initial matter power spectrum) and CDM dynamics (i.e., without accounting for wave-like evolution). While there is evidence from simulations that this is a reasonable approximation (Zhang et al., 2018; May & Springel, 2023), there are a number of other approaches for calculating the HMF in FDM. For instance, semi-analytic predictions of the HMF that also assume FDM IC and CDM dynamics (see e.g. Kulkarni & Ostriker 2021). Alternatively, one can attempt to model dynamical FDM effects. Even in linear theory, FDM predicts scale-dependent growth that differs from CDM growth on small scales (Marsh & Silk, 2013; Hlozek et al., 2015; Suárez & Chavanis, 2015; Li et al., 2019; Laguë et al., 2021), while spherical collapse may be further modified (Sreenath, 2019; Harko, 2019; Kurinchi-Vendhan et al., 2022). In fact, some semi-analytic models approximately account for some of these effects (Marsh & Silk, 2013; Marsh, 2016b; Du et al., 2017). On the other hand, the results of recent relatively large wave-dynamical simulations favor HMFs fairly similar to Schive et al. (2016) (May & Springel, 2023), although these simulations currently model FDM masses somewhat below $m_{22}=1$. The larger de Broglie wavelengths at smaller FDM masses make these cases easier to simulate, yet they lie beneath the range most relevant for our current constraints.
@@
First, we determine our conclusions with FDM ICs following the sharp $k$-space procedure of Kulkarni & Ostriker (2021). Then we consider FDM ICs plus dynamical suppression, approximated as an effective scale-dependent collapse threshold, $\delta_{c}(M,z)$, following Marsh & Silk (2013); Marsh (2016b).
Figure 5: Ratio of FDM to CDM UVLFs under different scenarios. FDM is harder to distinguish from CDM if observing at brighter luminosities or if the FDM particle is heavier. The deviation at higher redshift may be slightly larger (black, orange, and blue curves) or significantly larger (red and green curves) depending on the behavior of the FDM mass function model, which is currently uncertain.
Figure 5 shows the impact of using either of these alternative HMF formulations. The $y$-axis shows the ratios of the FDM to CDM UVLF. [^10] For the sharp-$k$ space curve the comparison is made to the best-fit UVLF when the CDM HMF is also calculated using a sharp-$k$ space smoothing window. At a given observing magnitude and $m_{22}$ the predictions using the Schive et al. (2016) HMF (black), sharp-$k$ space HMF (Section 5.1) (red) and (Marsh, 2016b) HMF (Section 5.2) (green) are similar for $z\lesssim 10$. However, at higher redshifts (which will be probed with JWST) these alternatives predict greater suppression in the UVLF. The yellow and blue curves respectively show that FDM becomes harder to distinguish from CDM for larger $m_{22}$ or when observing only brighter galaxies.
@@ 5.1 Sharp-$k$ Window Halo Mass Function
A naive calculation of $\sigma_{M}$ using the FDM initial power spectrum filtered with a sharp spherical smoothing window of radius $R=\left[3M/4\pi\bar{\rho}_{m}\right]^{1/3}$ (i.e. a real-space top-hat) predicts the existence of unphysical halos below the Jeans scale (Benson et al., 2012; Schneider et al., 2013; Ureña-López & Gonzalez-Morales, 2016; Leo et al., 2018; Kulkarni & Ostriker, 2021). This results because averaging over a top-hat window in real space (or indeed a gaussian window $\tilde{W}(kR)\propto e^{-(kR)^{2}}$ (Schneider et al., 2013)) not only includes the relevant scales but also includes a contribution from re-weighting segments of longer wavelength fluctuations (Benson et al., 2012; Schneider et al., 2013; Kulkarni & Ostriker, 2021). This is a negligible effect in CDM but it can dominate at small scales for theories like FDM or WDM which predict a sharp truncation in power at those scales.
An alternative choice is a sharp-$k$ window function (a top-hat in $k$-space) (Bond et al., 1991). [^11] Another possibility is a smooth $k$-space filter (Leo et al., 2018). This filter produces similar results to the real space top-hat in CDM (Schneider et al., 2013), and it explicitly excises unwanted longer wavelength (smaller $k$) modes when calculating $\sigma_{M}$. One condition is this filter requires an appropriate assignment of real space scale, $R\sim 2.5/k$ (Benson et al., 2012; Kulkarni & Ostriker, 2021), and rescaling of $\delta_{c}\mapsto 1.195\delta_{c}$ (Kulkarni & Ostriker, 2021). Using this filter with an FDM initial matter power spectrum predicts a much sharper suppression in the HMF compared to Schive et al. (2016). This suppression still onsets near $M_{1/2}$ at $z\sim 5-10$ although it shifts to slightly more massive halos at early times. The red curve in Figure 5 illustrates that this model gives larger differences in the UVLF (between FDM and CDM) at high redshift than in the Schive et al. (2016) formula, at least for the example FDM mass shown.
@@
The difference in behavior may be due to this simple semi-analytic model missing important nonlinear effects (Kulkarni & Ostriker, 2021). Alternatively, the Schive et al. (2016) simulations may suffer from simulation artifacts (such as unremoved “spurious halos” (Wang & White, 2007; Schive et al., 2016)). In addition, the fitting formula itself may not extrapolate well outside its calibration range in particle mass ($m_{22}\sim 1-3$), halo mass $M\gtrsim 10^{9}M_{\odot}$, or possibly redshift (Schive et al., 2016). Yet, current large FDM wave simulations (May & Springel, 2023) do not indicate as sharp a cutoff as predicted by the sharp-$k$ mass function.
@@ 5.2 Scale-Dependent Collapse and Quantum Pressure
Besides the effect on the initial matter distribution, the dynamical consequences of the uncertainty principle may be important. The zero-point momenta of FDM particles implies a force which opposes gravitational confinement, known as quantum pressure121212or sometimes quantum potential; it appears as an explicit term in the fluid formulation of the Schrödinger Equation (Madelung, 1926; Feynman, 1965). (QP), slowing growth and forbidding structure below the local Jeans scale (Marsh & Silk, 2013; Hui et al., 2017; Ferreira, 2021). This effect is thought to be subdominant and is so far neglected in our analysis, although its relative importance is still being investigated (Zhang et al., 2018; May & Springel, 2023).
One approximate HMF modeling approach, intended to roughly account for FDM dynamical effects, is to introduce a scale-dependent collapse threshold, $\delta_{c}(M,z)$ into the excursion-set based models (Marsh & Silk, 2013; Bozek et al., 2015; Marsh, 2016b; Du et al., 2017). If dynamical effects should hinder the formation of halos of mass $M$ at redshift $z$, a larger $\delta_{c}(M,z)$ leads to fewer of those kinds of halos. This can provide a sharper cutoff in the mass function, even when $\sigma_{M}$ is calculated using a real-space top-hat window function, and a complete suppression in halo formation below local Jeans scales.
For a specific example of this modeling approach, we consider the fitting formula from Marsh (2016b) [^13] There is a slight difference between the formula given in the text and the accompanying GitHub repository WarmAndFuzzy (Marsh, 2017). Including QP should lead to additional suppression. We therefore implement the GitHub version as the text version lies above (Schive et al., 2016) at $z<8$. of a mass-dependent barrier based on the model of Marsh & Silk (2013). In this case, $\delta_{c}(M,z)$ is 1.686 at high mass but exponentially increases near the suppression mass scale. Across $z=5-10$ this HMF matches Schive et al. (2016) well for $M\gtrsim M_{1/2}$ but is more sharply suppressed at smaller masses. At higher redshifts this mass function predicts an additional suppression at somewhat higher halo mass (note the green curve in Figure 5).
@@
However, the differences become stark when predicting the impact on future UVLF surveys, such as from JWST. While the Schive et al. (2016) fitting function evolves with redshift in the same way as CDM, the Marsh (2016b) HMF predicts halos near the cutoff mass will take much longer to collapse and so will be rarer at early times. If this principle manifests in the true FDM HMF, then JWST UVLF measurements should lead to much stricter bounds on the FDM particle mass (see Section 7). This motivates further study of FDM dynamical effects.

### v2

@@ 4 Sensitivity to Model Assumptions
The faint end of the UVLF, even including the faint, lensed galaxies of B22, is well-fit by the double power-law luminosity-halo mass relation (Sipple & Lidz, 2024). This assumption is also consistent with results from zoom-in hydrodynamic simulations of galaxy formation from the FIRE collaboration (Ma et al., 2018). Yet, it also may be the case that feedback effects (which arrest star formation) are weaker (e.g. if star formation is exceptionally “bursty" Furlanetto & Mirocha 2022; Sun et al. 2023) at the faint-end, or non-hierarchical structure formation in FDM may lead to younger, brighter stars in the lowest-mass halos (Corasaniti et al., 2017; Ni et al., 2019; Shen et al., 2024). [^8] The possibility of stronger feedback effects in the smallest halos (e.g. from radiative feedback during the reionization process, Borrow et al. 2023) is also not captured by this simple model. However, we do not find a clear turnover in the current UVLF data and, unlike the previous scenarios, neglecting this effect in our analyses leads to only a more conservative bound on FDM mass.
@@ 4.1 Cumulative UVLF
The strong suppression in the FDM halo mass function implies the cumulative number density of halos asymptotes to some finite value. This can be compared with the total observed abundance of galaxies above the limiting luminosity of the B22 data. Thus, we can provide a more conservative constraint by assuming that only one or fewer galaxies occupy each halo on average. This is expected theoretically at the faint end (Ma et al., 2018) and supported empirically by stringent bounds on satellite fractions at high redshift (which are only $\mathcal{O}\left(1\%\right)$ at $z\sim 5-6$ (Harikane et al., 2022), and may be even smaller at the lower luminosities and higher redshifts most relevant here). While our previous approach uses the full shape of the UVLF and delivers tighter bounds, the cumulative UVLF technique involves fewer assumptions regarding how galaxies populate their host halos. However, in the combined B21+B22 data set, the observed cumulative number density of halos is dominated by faint galaxies only detectable in the B22 lensed survey. For completeness we will also provide bounds using only the B21 field measurements which are more robust to systematic errors. Although the cumulative UVLF approach has been applied in previous work (Schive et al., 2016; Menci et al., 2017; Ni et al., 2019), we emphasize here its robustness to galaxy-halo connection modeling uncertainties.
@@
Given the requirement that fewer than one galaxy resides in each halo, on average, and with $M_{\rm min}$ denoting the minimum mass of dark matter halos hosting star formation, the halo mass function is inconsistent with the cumulative galaxy number density implied by UVLF measurements [^9] The total is driven almost entirely by the faintest resolved bins. Including blank-field B21 measurements does not improve our constraint. unless:
@@ 4.2 Modelling Enhanced Star Formation in Small Halos
If star formation proceeds differently in the smallest halos, we may still be able to model this alternate behavior as a simple additional power law in our UVLF modeling procedure. To this end, we introduce two additional degrees of freedom in our $L_{c}(M)$ relation, a new power law index, $p_{2}$, and its relevant mass scale, $M_{2}$, (similar to the “shallow-slope" model of Sipple & Lidz 2024). That is, we suppose $L_{c}(M)\rightarrow L_{c}(M)\times(M/M_{2})^{p_{2}-p}$ for $M\leq M_{2}$ and perform a new fit with a vector of parameters $\mathbf{\theta}_{2}=(p,q,r,M_{\rm UV,0},\log_{10}(M_{1}/\textrm{M}_{\odot}),1/m_{22},p_{2},\log_{10}(M_{2}/\textrm{M}_{\odot}))$ with the same priors on the original six parameters and conservatively chosen priors for $(p_{2},\log_{10}(M_{2}/\textrm{M}_{\odot}))$ of ([8, 14], [-3.5, 3.5]).
Sampling these eight parameters with MCMC, we find this model produces a similar constraint to the cumulative UVLF, (here $m_{22}>5.8$ at $2\sigma$) (Table 2). If the very smallest halos are actually bright enough to be included in B22 (necessitating 23$\sim$ 2-3$\times$ 2 - 3 × the predicted star formation efficiency of our fiducial model), this could somewhat compensate for the fewer halos at a given mass in FDM. While this may indicate more efficient star formation at the faint-end, including additional parameters always presents a danger of overfitting. This new behavior coincidentally onsets near the edge of observational capabilities, where the errors bars are relatively large. Further progress may require improving the statistical precision of the faint-end UVLF measurements and additional simulations exploring star formation in small mass dark matter halos.
In contrast, the existence of greater scatter in the $L(M)$ relation for low mass halos, potentially due to bursty star formation (i.e., a short star formation duty cycle), could strengthen the limit on FDM particle mass. If only a fraction of the $\sim 10^{9}-10^{10}M_{\odot}$ mass halos are UV bright enough to be detected in the B22 dataset, then some of the observed galaxies must necessarily reside in still smaller mass halos to match the observed galaxy abundance.
@@ 4.3 Effect of Prior
Another common choice in the literature is a flat prior in $\log_{10}(m_{\rm FDM})$ (Rogers & Peiris, 2021; Dentler et al., 2022; Winch et al., 2024), or in the log of the suppression halo mass scale (Nadler et al., 2021; Banik et al., 2021). This gives equal weight to different regions in log-space, but only over a finite interval, and so this cannot explicitly consider CDM where $m_{22}\rightarrow\infty$. The upper bounds on these intervals are chosen taking into account experimental sensitivity, although if the likelihood is not well-peaked in the prior region this choice can somewhat influence the confidence intervals on $m_{22}$. Here we choose an upper bound of $m_{22}=100$ ($M_{1/2}\approx 10^{8}\text{M}_{\odot}$) as the current UVLF data do not probe small enough halo masses to distinguish heavier FDM particle masses from CDM. With a prior of $\log_{10}(m_{22})\in[-1,2]$ (instead of our fiducial case of $1/m_{22}\in[0,2]$) our $2\sigma$ UVLF modeling constraint shifts only slightly from $m_{22}>15$ to $m_{22}>18$ (Tables 1,2).
@@ 5 FDM Halo Mass Function Modelling Uncertainties
Although a great deal of progress has been made recently, there is not yet a single simulation which spans the large dynamic range in scale need to resolve FDM effects near the local de Broglie wavelength, $\lesssim 100$ pc, while capturing statistical behavior on cosmological scales, $\gtrsim 100$ Mpc (Zhang et al., 2018a; Li et al., 2019; Hui, 2021; May & Springel, 2023; Laguë et al., 2024). The halo mass function relation of Equation 1 is only one approximation for how structure formation occurs in FDM. As mentioned in Section 2.1, it is fit to $N$-body simulations of Schive et al. (2016) with FDM initial conditions (IC) (including a suppressed initial matter power spectrum) and CDM dynamics (i.e., without accounting for wave-like evolution). While there is evidence from simulations that this is a reasonable approximation (Zhang et al., 2018b; May & Springel, 2023), there are a number of other approaches for calculating the HMF in FDM. For instance, semi-analytic predictions of the HMF that also assume FDM IC and CDM dynamics (see e.g. Kulkarni & Ostriker 2022). Alternatively, one can attempt to model dynamical FDM effects. Even in linear theory, FDM predicts scale-dependent growth that differs from CDM growth on small scales (Marsh & Silk, 2014; Hlozek et al., 2015; Suárez & Chavanis, 2015; Li et al., 2019; Laguë et al., 2021), while spherical collapse may be further modified (Sreenath, 2019; Harko, 2019; Kurinchi-Vendhan et al., 2022). In fact, some semi-analytic models approximately account for some of these effects (Marsh & Silk, 2014; Marsh, 2016a; Du et al., 2017). On the other hand, the results of recent relatively large wave-dynamical simulations favor HMFs fairly similar to Schive et al. (2016) (May & Springel, 2023), although these simulations currently model FDM masses somewhat below $m_{22}=1$. The larger de Broglie wavelengths at smaller FDM masses make these cases easier to simulate, yet they lie beneath the range most relevant for our current constraints.
@@
First, we determine our conclusions with FDM ICs following the sharp $k$-space procedure of Kulkarni & Ostriker (2022). Then we consider FDM ICs plus dynamical suppression, approximated as an effective scale-dependent collapse threshold, $\delta_{c}(M,z)$, following Marsh & Silk (2014); Marsh (2016a).
Figure 5: Top: Comparison of different FDM halo mass function models for $m_{22}=15$ at $z=8,15$. The solid curves show CDM predictions. The fiducial case is shown in black while the gray line uses a sharp $k$-space window function (see Section 5.1). The Schive, Sharp-$k$, and Marsh broken colored curves represent our fiducial HMF choice, fit to simulations, (Schive et al. 2016, Equation 1), an alternate analytic approach (Kulkarni & Ostriker 2022, Section 5.1), and one including the impact of dynamical suppression effects from quantum pressure (Marsh & Silk 2014; Marsh 2016a, Section 5.2, see Footnote 13) respectively. Bottom: The ratio of each FDM mass function to the corresponding CDM HMF.
Figure 5 compares these two alternate halo mass function approaches to the Schive et al. (2016) halo mass function for an FDM mass of $m_{22}=15$. The black solid curve shows the typical CDM prediction, which should be compared with the FDM predictions from the Schive et al. (2016) and Marsh (2016a) mass functions. The gray solid curve shows the CDM HMF calculated with the sharp-$k$ window function which should be compared to the FDM HMF calculated with the sharp-$k$ filter.
Figure 6: Ratio of FDM to CDM UVLF values in a particular luminosity bin across redshift under different scenarios. FDM is harder to distinguish from CDM if observing at brighter luminosities or if the FDM particle is heavier. The deviation at higher redshift may be slightly larger (black, orange, and blue curves) or significantly larger (red and green curves) depending on the behavior of the FDM mass function model, which is currently uncertain.
Figure 6 shows the impact of using either of these alternative HMF formulations. The $y$-axis shows the ratios of the FDM to the corresponding CDM UVLF. At a given observing magnitude and $m_{22}$ the predictions using the Schive et al. (2016) HMF (black), sharp-$k$ space HMF (Section 5.1) (red) and (Marsh, 2016a) HMF (Section 5.2) (green) are similar for $z\lesssim 10$. However, at higher redshifts (which will be probed with JWST) these alternatives predict greater suppression in the UVLF. The yellow and blue curves respectively show that FDM becomes harder to distinguish from CDM for larger $m_{22}$ or when observing only brighter galaxies.
@@ 5.1 Sharp-$k$ Window Halo Mass Function
A naive calculation of $\sigma_{M}$ using the FDM initial power spectrum filtered with a sharp spherical smoothing window of radius $R=\left[3M/4\pi\bar{\rho}_{m}\right]^{1/3}$ (i.e. a real-space top-hat) predicts the existence of unphysical halos below the Jeans scale (Benson et al., 2013; Schneider et al., 2013; Ureña-López & Gonzalez-Morales, 2016; Leo et al., 2018; Linares Cedeño et al., 2021; Kulkarni & Ostriker, 2022). This results because averaging over a top-hat window in real space (or indeed a gaussian window $\tilde{W}(kR)\propto e^{-(kR)^{2}}$ (Schneider et al., 2013)) not only includes the relevant scales but also includes a contribution from re-weighting segments of longer wavelength fluctuations (Benson et al., 2013; Schneider et al., 2013; Kulkarni & Ostriker, 2022). This is a negligible effect in CDM but it can dominate at small scales for theories like FDM or WDM which predict a sharp truncation in power at those scales.
An alternative choice is a sharp-$k$ window function (a top-hat in $k$-space) (Bond et al., 1991). [^11] Another possibility is a smooth $k$-space filter (Leo et al., 2018). This filter produces similar results to the real space top-hat in CDM (Schneider et al., 2013), and it explicitly excises unwanted longer wavelength (smaller $k$) modes when calculating $\sigma_{M}$. One condition is this filter requires an appropriate assignment of real space scale, $R\sim 2.5/k$ (Benson et al., 2013; Kulkarni & Ostriker, 2022), and rescaling of $\delta_{c}\mapsto 1.195\delta_{c}$ (Kulkarni & Ostriker, 2022). As shown in Figure 5, using this filter with an FDM initial matter power spectrum predicts a much sharper suppression in the HMF compared to Schive et al. (2016). This suppression still onsets near $M_{1/2}$ and has a weak dependence on redshift. Still, the red curve in Figure 6 illustrates that this model can sometimes lead to larger differences in the UVLF (between FDM and CDM) at high redshift than in the Schive et al. (2016) scenario. The redshift dependence of $L_{c}(M,z)$ (Equation 2, Table 1) implies that the same luminosity bin at higher redshift probes smaller halos. Around the FDM suppression scale, this difference is more dramatic in the sharp-$k$ case.
@@
The difference in behavior may be due to this simple semi-analytic model missing important nonlinear effects (Kulkarni & Ostriker, 2022). Alternatively, the Schive et al. (2016) simulations may suffer from simulation artifacts (such as unremoved “spurious halos" (Wang & White, 2007; Schive et al., 2016)). In addition, the fitting formula itself may not extrapolate well outside its calibration range in particle mass ($m_{22}\sim 1-3$), halo mass $M\gtrsim 10^{9}M_{\odot}$, or possibly redshift (Schive et al., 2016). Yet, current large FDM wave simulations (May & Springel, 2023) do not indicate as sharp a cutoff as predicted by the sharp-$k$ mass function.
@@ 5.2 Scale-Dependent Collapse and Quantum Pressure
Besides the effect on the initial matter distribution, the dynamical consequences of the uncertainty principle may be important. The zero-point momenta of FDM particles implies a force which opposes gravitational confinement, known as quantum pressure121212or sometimes quantum potential; it appears as an explicit term in the fluid formulation of the Schrödinger Equation (Madelung, 1926; Feynman, 1965). (QP), slowing growth and forbidding structure below the local Jeans scale (Marsh & Silk, 2014; Hui et al., 2017; Ferreira, 2021). This effect is thought to be subdominant and is so far neglected in our analysis, although its relative importance is still being investigated (Zhang et al., 2018b; May & Springel, 2023).
One approximate HMF modeling approach, intended to roughly account for FDM dynamical effects, is to introduce a scale-dependent collapse threshold, $\delta_{c}(M,z)$ into the excursion-set based models (Marsh & Silk, 2014; Bozek et al., 2015; Marsh, 2016a; Du et al., 2017). If dynamical effects should hinder the formation of halos of mass $M$ at redshift $z$, a larger $\delta_{c}(M,z)$ leads to fewer of those kinds of halos. This can provide a sharper cutoff in the mass function, even when $\sigma_{M}$ is calculated using a real-space top-hat window function, and a complete suppression in halo formation below local Jeans scales.
For a specific example of this modeling approach, we consider the fitting formula from Marsh (2016a) [^13] There is a slight difference between the formula given in the text and the accompanying GitHub repository WarmAndFuzzy (Marsh, 2017). Including QP should lead to additional suppression. We therefore implement the GitHub version as the text version lies above (Schive et al., 2016) at $z<8$. of a mass-dependent barrier based on the model of Marsh & Silk (2014). In this case, $\delta_{c}(M,z)$ is 1.686 at high mass but exponentially increases near the suppression mass scale. Across $z=5-10$ this HMF matches Schive et al. (2016) well for $M\gtrsim M_{1/2}$ but is more sharply suppressed for much smaller masses. As shown in Figure 5, at higher redshifts this mass function predicts an additional suppression at somewhat higher halo mass (also note the green curve in Figure 6).
@@
However, the differences become stark when predicting the impact on future UVLF surveys, such as from JWST. While the Schive et al. (2016) fitting function evolves with redshift in the same way as CDM, the Marsh (2016a) HMF predicts halos near the cutoff mass will take much longer to collapse and so will be rarer at early times. If this principle manifests in the true FDM HMF, then JWST UVLF measurements should lead to much stricter bounds on the FDM particle mass (see Section 7). This motivates further study of FDM dynamical effects.